# Make an Offer They Can’t Refuse: Grounding Bayesian Persuasion in Real-World Dialogues without Pre-Commitment

### v1

@@ 1. Introduction
Persuasion is a fundamental form of human social interaction, enabling individuals to influence others’ beliefs and decisions through communication brembeck_persuasion_1976 . While large language models (LLMs) already exhibit strong abilities in language generation and understanding, they remain limited in strategic persuasion tasks: LLMs often fail to effectively exploit information asymmetry and struggle to design messages that rationally shift a persuadee’s beliefs.
Bayesian Persuasion (BP), a game-theoretic framework for information design, offers a complete solution in constrained mathematical settings by modeling how a persuader can disclose information to maximize desired actions kamenica_bayesian_2011 . However, applying BP in open-ended natural language dialogue raises a key challenge: the mathematical constructs of BP—such as priors, world states, and posterior updates—must be verbalized into coherent and persuasive arguments. For example, when convincing friends to try a new restaurant, the priors (the probability that the restaurant will suit their taste) and utilities (how much they would enjoy it) are rarely stated explicitly. The offer would be refused if the story is not that convincing. Moreover, a pivotal step in BP is the sender’s commitment to a signaling schema, which we argue is crucial for natural language realization.
Existing approaches, such as the method in li_verbalized_2025 , typically rely on pre-commitment by statically encoding the schema in the persuadee’s prompt, bypassing the need for its communication. Consequently, there is a lack of systematic methods for implementing BP where the schema itself is articulated within a single-turn natural language interaction under information asymmetry.
@@
Motivated by preliminary observations that powerful LLMs such as GPT-5 openai_introducing_2025 and DeepSeek-V3.1 deepseek-v31_2025 display BP-like reasoning capability when prompted with above mechanism carefully, we investigate two key questions: (1) Can LLMs consistently execute BP strategies through natural language to enhance persuasion success? And (2) can supervised fine-tuning empower smaller models to achieve BP performance comparable to larger counterparts?
@@ 2.1. Persuasion with Large Language Models
Recent research has increasingly explored the persuasive capabilities of LLMs, examining both their potential benefits and societal risks. ramani_persuasion_2024 investigate multi-agent frameworks that could enhance persuasion efficacy through collaborative specialization, where auxiliary agents handle tasks like strategy development and resistance analysis. gemp_steering_2024 introduce equilibrium solvers that guide LLM dialogue generation, enabling models to reason about consistent strategic interactions. shi_persuasion_2025 reviews AI persuasion systems from a social good perspective, offering theoretical support for discussing ethical implications. cheng_towards_2025 develop theory-driven evaluation frameworks to systematically measure LLMs’ ability to change beliefs and decisions.
Additionally, several studies are focused on building specialized datasets. Persuasion for Good wang_persuasion_2020 examines how individual traits affect persuasion outcomes and provides a basis for strategy adaptation. jin-etal-2024-persuading ; hayati-etal-2020-inspired build datasets for the persuasion in daily scenarios. CToMPersu zhang_persuasion_2025 proposes a “double-blind” framework where persuasion strategies and mental states remain undisclosed, which is particularly well-suited for BP analysis due to its explicit information asymmetry design.
@@ 2.2. Bayesian Persuasion and Information Design
As an information design method, BP provides a normative framework for understanding how an informed sender can design signals to influence a receiver’s beliefs and actions. The seminal work of kamenica_bayesian_2011 characterizes when persuasive signaling benefits the sender and derives optimal signaling schemes. This foundation has been expanded through the broader lens of information design bergemann_information_2016 ; bergemann_information_2019 , which unifies BP with communication in games and robust prediction.
Algorithmic perspectives on information design are comprehensively reviewed by dughmi_algorithmic_2017 . Some studies adapt BP to dynamic and uncertain environments. castiglioni_online_2020 ; bernasconi_optimal_2023 ; bacchiocchi_online_2024 address online extensions of BP and continuously improve the regret bound. bacchiocchi_without_2024 tackle the fully prior-free case. Other directions include persuasion with externalities where one receiver’s actions affect others’ utilities shaki_bayesian_2025 and complexity analyses showing that discovering persuasive messages is NP-hard while verifying them is tractable wojtowicz_when_2024 .
Beyond theoretical extensions, BP has been applied to practical AI challenges. bai_efficient_2024 employ BP for model-agnostic alignment, using a smaller model as an advisor that sends signals to guide larger models’ responses, demonstrating BP’s utility in improving model performance without direct training. zhang2025roadmap introduce BP into AI alignment and governance, proposing that AI can be made to act according to human intentions by designing information during the post-deployment phase.
@@ 2.3. Verbalizing Bayesian Persuasion in Natural Language
Translating BP into natural language requires verbalizing signals, beliefs, and posterior updates within coherent narratives. li_verbalized_2025 address this through Verbalized Bayesian Persuasion (VBP), which incorporates the signaling scheme directly into the receiver’s prompt and adjusts sender generation via keyword manipulation. While effective in their evaluated domains, this approach relies on explicit prompt engineering and is demonstrated in limited, scenario-specific settings.
@@ 4.1.1. Dataset and Bayesian-Setup Construction
We build our experimental corpus from the CToMPersu dataset zhang_persuasion_2025 , which provides multi-domain persuasion scenarios with explicit information asymmetry design. Each case includes a persuader, a persuadee with theory of mind annotations carruthers1996theories , a background story, and a persuasion goal. To ground these scenarios in the BP framework, we augment each case with a structured Bayesian setup. This setup specifies a two-state world (Positive/Negative), prior beliefs, receiver actions (Accept/Reject), fixed sender utilities ({1, 0}), and state-dependent receiver utilities. Crucially, we ensure that under the no-information setting, the receiver’s expected utility is slightly lower than zero, creating natural resistance that must be overcome through strategic information disclosure. An example is shown in the Appendix A.1. These configurations were automatically generated using DeepSeek-V3 deepseekai2024deepseekv3technicalreport and validated for consistency.
@@ 4.1.2. Settings and Conditions
Table 1. Experimental settings overview. Abbreviations: SD = Self-derived, SCE = Scenario, NUM = Bayesian setup, UTL = Verbalized prior & receiver utility, DEF = BP definitions, VER = Verbalization prompt, SMP = Self-modeling prompt, RAT = Rational prompt. NBP baselines include Naive and Strong for Explicit view (with specific methods like Logical appeal wang_persuasion_2020 ), and Strong for Self-derived view (with alternative strategic principles such as Nash equilibrium Nash_1950 ).
@@
  • Strong: Enhanced NBP persuasion with game-theoretic principles or appeal methods inspired by wang_persuasion_2020 .
@@ 4.1.3. Models and Evaluation Protocol
We evaluate eight persuaders across diverse capabilities: DeepSeek-V3.1111Unless stated, DeepSeek-V3.1 refers to thinking/reasoning mode deepseek-v31_2025 , GPT-5 openai_introducing_2025 , Qwen3-4B yang2025qwen3technicalreport , Qwen3-0.6B yang2025qwen3technicalreport , and their supervised fine-tuned variants (Qwen3-4B, Qwen3-0.6B), plus Gemma-3-4B-it gemma_2025 and Gemma-3-1B-it gemma_2025 . [^2] To conserve space in the table, models are referred to by abbreviated names: V3.1, Qwen0.6B, Qwen4B, Gemma1B, and Gemma4B. Five models marked with also serve as persuadees. We run all $8\times 5$ persuader-persuadee pairs for each setting and competence pairing.
@@ 4.5.1. Evaluation Design and Procedure
For each comparison, participants rated five dimensions on a forced-choice basis, adapted from classical rhetorical analysis hidey_analyzing_2017 : Persuasiveness, Emotional Resonance, Credibility, Logical Coherence, and Fluency. Detailed definitions of these five evaluation dimensions are provided in Appendix A.5.2.
@@ 4.5.2. Human Evaluation Results and Analysis
Interestingly, comparison with LLM-as-a-judge evaluations (Table 10 and Figure 4 right panel) reveals both alignments and divergences in assessment criteria. Specifically, we employ DeepSeek-V3.2-Exp (Thinking & non-Thinking mode) deepseekai2024deepseekv32 , GPT-5, Qwen3-MAX yang2025qwen3technicalreport and Qwen3-235B-A22B-2507 yang2025qwen3technicalreport as the judge models. While humans prioritize emotional resonance and overall persuasiveness, LLM judges place greater emphasis on logical coherence and fluency, particularly favoring SFNL in logical dimension (55 preferences) compared to human evaluation (53). This discrepancy highlights the importance of multi-perspective evaluation in persuasive communication.

### v2

@@ 1. Introduction
Persuasion is a fundamental form of human social interaction, enabling individuals to influence others’ beliefs and decisions through communication Brembeck (1976). While large language models (LLMs) already exhibit strong abilities in language generation and understanding, they remain limited in strategic persuasion tasks: LLMs often fail to effectively exploit information asymmetry and struggle to design messages that rationally shift a persuadee’s beliefs.
Bayesian Persuasion (BP), a game-theoretic framework for information design, offers a complete solution in constrained mathematical settings by modeling how a persuader can disclose information to maximize desired actions Kamenica and Gentzkow (2011). However, applying BP in open-ended natural language dialogue raises a key challenge: the mathematical constructs of BP—such as priors, world states, and posterior updates—must be verbalized into coherent and persuasive arguments. For example, when convincing friends to try a new restaurant, the priors (the probability that the restaurant will suit their taste) and utilities (how much they would enjoy it) are rarely stated explicitly. The offer would be refused if the story is not that convincing. Moreover, a pivotal step in BP is the sender’s commitment to a signaling schema, which we argue is crucial for natural language realization.
Existing approaches, such as the method in Li et al. (2025), typically rely on pre-commitment by statically encoding the schema in the persuadee’s prompt, bypassing the need for its communication. Consequently, there is a lack of systematic methods for implementing BP where the schema itself is articulated within a single-turn natural language interaction under information asymmetry.
@@
Motivated by preliminary observations that powerful LLMs such as GPT-5 OpenAI (2025) and DeepSeek-V3.1 DeepSeek-AI (2025a) display BP-like reasoning capability when prompted with above mechanism carefully, we investigate two key questions: (1) Can LLMs consistently execute BP strategies through natural language to enhance persuasion success? And (2) can supervised fine-tuning empower smaller models to achieve BP performance comparable to larger counterparts?
@@ 2.1. Persuasion with Large Language Models
Recent research has increasingly explored the persuasive capabilities of LLMs, examining both their potential benefits and societal risks. Ramani et al. (2024) investigate multi-agent frameworks that could enhance persuasion efficacy through collaborative specialization, where auxiliary agents handle tasks like strategy development and resistance analysis. Gemp et al. (2024) introduce equilibrium solvers that guide LLM dialogue generation, enabling models to reason about consistent strategic interactions. Shi (2025) reviews AI persuasion systems from a social good perspective, offering theoretical support for discussing ethical implications. Cheng and You (2025) develop theory-driven evaluation frameworks to systematically measure LLMs’ ability to change beliefs and decisions.
Additionally, several studies are focused on building specialized datasets. Persuasion for Good Wang et al. (2020) examines how individual traits affect persuasion outcomes and provides a basis for strategy adaptation. Jin et al. (2024); Hayati et al. (2020) build datasets for the persuasion in daily scenarios. CToMPersu Zhang and Zhou (2025) proposes a “double-blind” framework where persuasion strategies and mental states remain undisclosed, which is particularly well-suited for BP analysis due to its explicit information asymmetry design.
@@ 2.2. Bayesian Persuasion and Information Design
As an information design method, BP provides a normative framework for understanding how an informed sender can design signals to influence a receiver’s beliefs and actions. The seminal work of Kamenica and Gentzkow (2011) characterizes when persuasive signaling benefits the sender and derives optimal signaling schemes. This foundation has been expanded through the broader lens of information design Bergemann and Morris (2016, 2019), which unifies BP with communication in games and robust prediction.
Algorithmic perspectives on information design are comprehensively reviewed by Dughmi (2017). Some studies adapt BP to dynamic and uncertain environments. Castiglioni et al. (2020); Bernasconi et al. (2023); Bacchiocchi et al. (2024a) address online extensions of BP and continuously improve the regret bound. Bacchiocchi et al. (2024b) tackle the fully prior-free case. Other directions include persuasion with externalities where one receiver’s actions affect others’ utilities Shaki et al. (2025) and complexity analyses showing that discovering persuasive messages is NP-hard while verifying them is tractable Wojtowicz (2024).
Beyond theoretical extensions, BP has been applied to practical AI challenges. Bai et al. (2024) employ BP for model-agnostic alignment, using a smaller model as an advisor that sends signals to guide larger models’ responses, demonstrating BP’s utility in improving model performance without direct training. Zhang et al. (2025) introduce BP into AI alignment and governance, proposing that AI can be made to act according to human intentions by designing information during the post-deployment phase.
@@ 2.3. Verbalizing Bayesian Persuasion in Natural Language
Translating BP into natural language requires verbalizing signals, beliefs, and posterior updates within coherent narratives. Li et al. (2025) address this through Verbalized Bayesian Persuasion (VBP), which incorporates the signaling scheme directly into the receiver’s prompt and adjusts sender generation via keyword manipulation. While effective in their evaluated domains, this approach relies on explicit prompt engineering and is demonstrated in limited, scenario-specific settings.
@@ 4.1.1. Dataset and Bayesian-Setup Construction
We build our experimental corpus from the CToMPersu dataset Zhang and Zhou (2025), which provides multi-domain persuasion scenarios with explicit information asymmetry design. Each case includes a persuader, a persuadee with theory of mind annotations Carruthers and Smith (1996), a background story, and a persuasion goal. To ground these scenarios in the BP framework, we augment each case with a structured Bayesian setup. This setup specifies a two-state world (Positive/Negative), prior beliefs, receiver actions (Accept/Reject), fixed sender utilities ({1, 0}), and state-dependent receiver utilities. Crucially, we ensure that under the no-information setting, the receiver’s expected utility is slightly lower than zero, creating natural resistance that must be overcome through strategic information disclosure. An example is shown in the Appendix A.1. These configurations were automatically generated using DeepSeek-V3 DeepSeek-AI (2024) and validated for consistency.
@@ 4.1.2. Settings and Conditions
Table 1. Experimental settings overview. Abbreviations: SD = Self-derived, SCE = Scenario, NUM = Bayesian setup, UTL = Verbalized prior & receiver utility, DEF = BP definitions, VER = Verbalization prompt, SMP = Self-modeling prompt, RAT = Rational prompt. NBP baselines include Naive and Strong for Explicit view (with specific methods like Logical appeal Wang et al. (2020)), and Strong for Self-derived view (with alternative strategic principles such as Nash equilibrium Nash (1950)).
@@
  • Strong: Enhanced NBP persuasion with game-theoretic principles or appeal methods inspired by Wang et al. (2020).
@@ 4.1.3. Models and Evaluation Protocol
We evaluate eight persuaders across diverse capabilities: DeepSeek-V3.1111Unless stated, DeepSeek-V3.1 refers to thinking/reasoning mode DeepSeek-AI (2025a), GPT-5 OpenAI (2025), Qwen3-4B Yang et al. (2025), Qwen3-0.6B Yang et al. (2025), and their supervised fine-tuned variants (Qwen3-4B, Qwen3-0.6B), plus Gemma-3-4B-it Team (2025) and Gemma-3-1B-it Team (2025). [^2] To conserve space in the table, models are referred to by abbreviated names: V3.1, Qwen0.6B, Qwen4B, Gemma1B, and Gemma4B. Five models marked with also serve as persuadees. We run all $8\times 5$ persuader-persuadee pairs for each setting and competence pairing.
@@ 4.5.1. Evaluation Design and Procedure
For each comparison, participants rated five dimensions on a forced-choice basis, adapted from classical rhetorical analysis Hidey et al. (2017): Persuasiveness, Emotional Resonance, Credibility, Logical Coherence, and Fluency. Detailed definitions of these five evaluation dimensions are provided in Appendix A.5.2.
@@ 4.5.2. Human Evaluation Results and Analysis
Interestingly, comparison with LLM-as-a-judge evaluations (Table 10 and Figure 4 right panel) reveals both alignments and divergences in assessment criteria. Specifically, we employ DeepSeek-V3.2-Exp (Thinking & non-Thinking mode) DeepSeek-AI (2025b), GPT-5, Qwen3-MAX Yang et al. (2025) and Qwen3-235B-A22B-2507 Yang et al. (2025) as the judge models. While humans prioritize emotional resonance and overall persuasiveness, LLM judges place greater emphasis on logical coherence and fluency, particularly favoring SFNL in logical dimension (55 preferences) compared to human evaluation (53). This discrepancy highlights the importance of multi-perspective evaluation in persuasive communication.